OpenStudy (anonymous):

antonia is making a cake she uses currants:walnuts:sultanas =4:3:5 mass=3.6 kg calculate mass of sultanas pls help

hartnn (hartnn):

since the ratio is 4:3:5 let the individual quantities be 4x, 3x and 5x total = 3.6 so, 4x+3x+5x = 3.6 can you find 'x' from here ? then sultanas' mass will be just 5x :)

OpenStudy (anonymous):

thank u

hartnn (hartnn):

welcome ^_^ if you want me to verify your answer, tell me what u got :)

OpenStudy (anonymous):

ya sure

OpenStudy (anonymous):

sorry , but i am getting the wrong answer pls help me

hartnn (hartnn):

sure, 4x+3x+5x = 3.6 (4+3+5)x =3.6 12x =3.6 x = 3.6/12 = 0.3 did you get the same 'x' value ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok now i got it

OpenStudy (anonymous):

thank u so much

hartnn (hartnn):

most welcome ^_^

OpenStudy (anonymous):

price=$18 given discount=15% work out the discount

hartnn (hartnn):

x% = x/100 so, 15% of 18 = 15/100 *18 =....?

OpenStudy (anonymous):

can i ask u if i come through more doubts

hartnn (hartnn):

sure! :) always happy to help ^_^

OpenStudy (anonymous):

7.15 hours to hours and minutes

hartnn (hartnn):

7.15 hours = 7 hours and 0.15 hours now we need to convert 0.15 hours to minutes x hrs = x*60 minutes 0.15 hrs = 0.15*60 minutes = ... ?

OpenStudy (anonymous):

wht does tht star mean

hartnn (hartnn):

multiplication

OpenStudy (anonymous):

k

OpenStudy (anonymous):

the mass of earth is app. 5.97*10 raise to 24 mass of moon is app. 1 upon 81 calculate mass of moon